The Life and Times of Studio ‘S’

Dicky Howett delves into the engineering logs of BBCtv Southampton.


Recently I’ve acquired a small set of BBC Television Engineering Logs written by BBC staff over short periods between 1986 and 1987. In this instance the logs refer to the BBC’s Southampton studio centre called with faultless logic, Studio ‘S’. As a rule, BBC engineering logs were used to chart the day-to-day progress (or lack of) concerning such technical trivia as loose cables, blown cue bulbs and indeed blown entire cameras!

Studio ‘S’ was (and probably still is) the studio for whence emanated such regional BBCtv delights as ‘South Today’. These BBC engineering logs (usually confidential-but not here) make interesting, pithy and occasionally arcane reading. Written by anonymous engineers, (they know who they are) the technical logs present an invaluable guide (where legible) of daily broadcast television engineering practices, plus trials and indeed many tribulations.

To set the scene, during the 1980s, Studio ‘S’ was a humble BBC television regional outpost, equipped with a hand-me-down collection of recalcitrant colour kit, incorporating four sturdy examples of that much-admired camera, the EMI 2001. However, as these logs graphically illustrate, by the end of the 1980s any residual love for the 2001 camera had completely ebbed away. The reasons for this (and more) are recorded thus………

STUDIO ‘S’ ENGINEERING LOG. (Abridged for intelligibility)

26/8/86.
CAM 2 Ghosting. Cable to viewfinder is faulty.
CAM 1 Zoom control is not centralised ie; is more sensitive one way than the other. Also tends to creep..

29/8/86.
CAM 4 Soft again this morning-are focus volts drifting? At lunchtime however the camera is now meshing. To cure meshing I re-adjusted focus to optimum and re-registered. Now OK.

4/9/86.
CAM 3 ‘Hooking’ slightly at top of picture. It occurs on all cameras but is worse for CAM 3

8/9/86.
CAM 1. Ring on Y channel. Found yolk earth strap not connected to yolk.

15/9/86.
CAM 1. Viewfinder low brightness. Swapped with MAINT spare.
CAM 1. Altho CAM 1 greyscales ok, on monitor it looks yellower than 2&3.

17/9/86.
ALL CAMERAS. An amazing increase in resolution and sensitivity achieved by cleaning the Autocue screens!

18/9/86.
PEDS. Tilt locks on 1 & 3 need tightening. Also pan lock on 2

19/9/86
CAM 3. 100Hz-ish pickup, mostly on Red, but slightly on other channels. Not visible with fully mod pix, so they are using it for now…

23/9/86.
ALL COLOUR MONITORS. Colour balance appalling!

30/9/86.
CAM 2. Excessive ringing on Green channel. Changed head amp. This improved matters.
CAM 1. Green tube. Small hole visible at low light levels.

2/10/86.
CAM 1. Green tube rejected for softness. Replaced with new tube. Okay.
CAM 1. Luminance meshing. Started off slight, increased dramatically, but eventually put right.

3/10/86.
CAM 2. Lens pack PSU S/C blowing main fuses.

8/10/86.
CAM 2. Nasty slash on Y channel. Cured by tightening earth tag on tube o/p.
CAM 1. Flashing during TX. (transmission)

9/10/86.
CAMS 1, 2 & 3. All cameras have had ringing and noisy channels.

13/10/86.
TUBE CONDITIONER TURNED ON FOR TEN HOURS.
Adjusted alignment of CAM 4. A little improvement.

29/10/86.
CAM 1. V/f has bogey top left. To be investigated.

30/10/86.
CAM 2. Slightly soft v/f. Duff peaking control. Replaced.
CAM 4. Soft due to duff zoom tracking.

6/11/86.
CAM 2. Zoom sticking.

10/11/86.
CAM2. Pedestal cable clamp shattered. (“It came off in my hand, guv..”) Quite..

23/11/86.
CAM 1. Meshing badly. Tweaked alignment so that meshing now matches that of CAM 2.

24/11/86.
Slide file slides (production) mechanically bust. Screw loose. Had my attention and sympathy. Now OK. Signed (DMB). (another hand had added “God help it!)

12/1/87.
CAM 1. Ped gassed up.

20/1/87.
CAM 2. Zoom tracking control changed due to existing one being bent Also, crap cleaned off front element of lens.

11/2/87.
CAM 1. Pan lock handle came loose and then came off in spigot which disappeared into its hole, with only a small bit showing. Tried unsuccessfully during line-up to retrieve spigot. Cameraman was happy to have no locking facility so left until tomorrow.

12/2/87.
CAM 1. Retrieved spigot from hole! Took front plate off pan/tilt head. Replaced locking handle and now OK.
CAM 2. Taken upstairs for major surgery. Block replaced with lovingly cleaned spare. Re-aligned, registered etc. No improvement noticeable. Bad ringing on all colour channels.

13/2/87.
CAM 2. Earth missing on Blue tube. Put in new earth tag and this cured ringing on all colour channels. Increased patch size. Checked beam align, focus. Still no vast improvement. Changed lens between CAM 1 & CAM 2.
CAM 1. Creeping zoom. Adjusted pot. Now OK.

23/2/87.
CAM 2 & 3. Both found with matrix out.
CAM 2. Blue flashing fixed.

24/2/87.
CAM 1. Very bad line rate shading on Red channel.

25/2/87.
CAM 1. (09.00) Shading on Red channel caused by viewfinder. Changed for spare.
CAM 1. (14.00). Camera died completely. Caused by new viewfinder.

4/3/87.
(18.45). All cameras looking naff. CAMS 1 & 2 not matching in colour.
CAM 3 sat up.

6/3/87.
CAM 1. Gassed up.
ALL CAMS. Yes, they do need a CCU line-up.

18/3/96.
CAM 4. Was seen to lose frame lock. Investigated but nothing concrete found.

23/3/87.
CAM 3. Pan & Tilt head locks need attention.

27/3/87.
CAM 4. Grey scale. Magenta cast now. Corrected. Gamma Laws don’t match. Needs a CCU line-up.
CAM 3. Hair-like object in top RH corner.

1/4/87.
CAM 1. April Fools Day!!. Out for shading corrector mod.
CAM 2. Field blanking not working. Traced fault back to shading corrector. Also failure on pulse board of field blanking o/p. Consistent line flashing possibly from camera end rather than CCU. Luminance yolk may have been jogged. Pictures softer than normal.

13/4/87.
CAM 2. (17.15ish) Viewfinder replaced with spare in an effort to reduce ringing on all channels.
CAM 2. (18.30). Replacement viewfinder causing breakthrough on picture with smells of burning. Replaced with original viewfinder but this also failed to work.

2/8/87.
CAM 1. Picture fluttering as last week (but not too serious).

As the writer of the log indicates ironically, however bad things may seem, the day is never completely lost with BBC Engineering on hand. Region television always had to make do with ‘roll-in’ cameras and second-hand kit., so the ground-rules were laid well in advance.

By the mid-1980s, the BBC’s main electronic production cameras were the Link 110 & 125 and on obs, the ubiquitous Philips LDK5. The staunch EMI 2001 soldiered on in various BBC backwaters, but by the end of the decade, the camera’s end was very much in sight. (It had originally been installed at TV centre Studio 8 in 1967 so its design life of ten years was well into extra time). Even so, the BBC doggedly maintained the camera at one major production centre. This was Studio ‘C’ at Elstree which until July 1991, recorded ‘EastEnders’. And it showed.

© Dicky Howett 2002